LPC meeting summary 07-11-2016 - final
Main purpose of the meeting: Discussion and status of the HI run commissioning and planning.
Jamie underlined the efforts made by the cryo team which managed to recover from the power outage due to the 400kV transformer intervention and thanked them for this.
He then mentioned the problems encountered over the weekend which delayed the commissioning. But in general the commissioning was still on schedule and first stable beam could in the best case expected on Tuesday morning around 4h.
For the LHCf fills it is considered that data could be taken as "end of fill" data taking, especially in case that the injectors would not be available for any reasons to prepare a new fill. The first ion re-fill is planned to happen during the RP survey in the SPS for the scheduled VIP visit. On a question from ALICE Rende answered that the survey needs to be done a certain time after the last beam in the SPS but he would check if some optimisation could be done.
While discussing the scheduling with the baseline criteria for ending the 5TeV part ALICE stressed that 600M min bias events would really be marginal for their physics programme and they would like to have a discussion before it will be decided that this threshold will be applied. Jamie remarked that no decision will be taken without discussion.
While discussing the VdM scans during the heavy ion5 TeV run Witold stated that it would be nice but (given the expected precision) not essential to have all scans in the same run. CMS underlined that they do not need this scan for physics but just do it to deliver some information for accounting. Witold will try to organise the scans with CMS. Jamie stated that experiments should organised among themselves these scan activities.
Jamie reminded that the LHCf run will be taken with separated beams in Point 1. He also underlined that care has to be taken to not let bunches which collide in IP 1/5/8 loose too much intensity so that they drop below the BPM sensitivities and therefore dump the beam which otherwise could still go on for a substantial time.
CMS (Christoph Roland) asked when this could happen. John answered that this would depend on the lifetimes of the beams which are not yet known. C.Roland remarked that CMS observed a rate of 200Hz in the commissioning fill and wanted to know if this can be improved. John remarked that there might be room for some improvement but not a lot. There might be a possibility to give more luminosity to IP 1/5 with different filling schemes, but this depends on how the run will evolve.
ALICE wanted to know when the VdM scans in the 8TeV part would be foreseen. Jamie answered that these details have not yet been planned and before planning them one would need to see how the 8 TeV run evolves. Witold underlined that some planning ahead would be necessary in order to have the availability of the experts (also on the machine side: BI) assured. Probably the scans will be done with 2 scans in one fill (i.e. all four scans in 2 fills). LPC will follow up with the BI experts.
LHCb plans to use the third fill with the production filling scheme to inject gas and start the SMOG measurements. This will distort the online lumi measurements of LHCb. Regarding the online lumi in the 8TeV part, LHCb will scale the 5TeV calibration to 8TeV.
Atlas cannot make use of the 2013 calibration since there luminosity detectors changed significantly. They will rely on the observed total cross section. According to Christoph Roland and Witold, CMS will do the same.